it tom covered the first 200 miles of a trip going 50mph, and the last 100 miles going 40mph, what was the average speed?
Answer by addingup(3677)

200+100= 300 Total trip was 300.
200/300= 2/3 of his trip at 50mph
100/300= 1/3 of his trip at 40mph
2/3*50+1/3*40= 2/3*50/1+1/3*40/1
= (2*50)/3*1)+(1*40)/(3*1)
= 100/3+40/3
= 33.333+13.333
= 46.667mph was his average speed.